Complete combustion of 7.30 g of a hydrocarbon produced 23.5 g of CO2 and 8.00 g of H2O. What is the empirical formula for the hydrocarbon? Insert subscripts as necessary.
Answer #1

Moles of CO2 = moles of C atoms in Hydrocarbon = mass / molar mass = 23.5 / 44 = 0.534 moles

moles of H2O = 0.5 * moles of H in the sample = mass / molar mass = 8 / 18 = 0.444 moles

moles of H atoms = 2 * 0.444 = 0.888 moles

ratio of moles of C to H = 0.534 / 0.888 = 3 / 5

Empirical formula = C3H5